OpenStudy (anonymous):

okie! there's a 6% commission on each car a dude sells. What's the commission on a car he sells for 7200?

OpenStudy (anonymous):

6/100*x=7200 0.06*x=7200 0.06x=7200 /0.06 /0.06 x=120000 the commission is 120000